2) For the function y = x2 - 4x - 5, perform the following tasks:
a) Put the function in the form y = a(x - h)2 + k.
Answer:
Show work in this space



b) What is the equation for the line of symmetry for the graph of this function?
Answer:



c) Graph the function using the equation in part a. Explain why it is not necessary to plot points to graph when using y = a (x – h)2 + k.
Show graph here.


Explanation of graphing.




d) In your own words, describe how this graph compares to the graph of y = x2?
Answer:

Answer by stanbon(75887)

2) For the function y = x2 - 4x - 5, perform the following tasks:
a) Put the function in the form y = a(x - h)2 + k.
Answer:
Show work in this space
x^2-4x +? = y+5
Complete the square on the left side and keep the equation balanced.
x^2-4x+4 = y+5+4
Factor to get:
(x-2)^2 = y+9
y=(x-2)^2 - 9


b) What is the equation for the line of symmetry for the graph of this function?
Answer:
EQUATION: x=2 is the line of symmetry.


c) Graph the function using the equation in part a. Explain why it is not necessary to plot points to graph when using y = a (x – h)2 + k.
Show graph here.

Explanation of graphing.
vertex is at (h,k); axis of symmetry is x=h; If a is positive the parabola
opens up; if a is negative the parabola opens down.



d) In your own words, describe how this graph compares to the graph of y = x2?
Answer:
y=(x-2)^2 - 9
Start with y=x^2
the (x-2) moves the graph two to the right
the -9 moves all points of the graph 9 down
---------------------
Cheers,
stan H.